# REPARO: Compositional 3D Assets Generation with Differentiable 3D Layout Alignment

### v1

@@ 4.2 Main results
Figure 5: User study of different models.
User study. We conducted a user study to compare our method with others. We gathered 280 responses from 40 human participants. Each participant was shown a reference image alongside four 3D assets (including our model and baseline model DreamGaussian [48], Wonder3D [32] and LRM [14]) simultaneously and asked to select the most realistic assets based on geometry, texture quality, and accurate placement. All options were presented in a randomized order with no time constraints. Figure 5 illustrates that our approach significantly outperformed previous methods in terms of human preference.
@@ 4.3 Ablation study
Table 2: Comparison of different loss functions in differentiable rendering.
(a) DreamGaussian-based REPARO.
$L_{a}$ $L_{s}$ CLIP score $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
0.833 17.296 0.826 0.233
0.809 17.820 0.849 0.210
0.833 17.391 0.827 0.232
(b) TripoSR-based REPARO.
$L_{a}$ $L_{s}$ CLIP score $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
0.822 17.765 0.865 0.216
0.813 17.906 0.867 0.212
0.823 17.778 0.865 0.216
Ablation for different loss functions in differentiable rendering. We conducted ablation experiments on different loss components in Eq. 9, with the results presented in Table 2. The results indicate that incorporating the high-level semantic loss function $L_{s}$ can bring improvements in both SSIM and perceptual LPIPS. However, the enhancement in CLIP score is not as pronounced. Additionally, when only $L_{s}$ is used, there is a notable decrease in CLIP score. For instance, in the DreamGaussian-based model, the CLIP score decreased by $2.1\%$. This suggests that relying solely on feature supervision can reduce the effectiveness of layout alignment.
Ablation for long-range appearance loss term. As shown in Eq. 5, our proposed long-range appearance loss $L_{a}$ incorporates RGB color distance, position distance, and depth value distance. More importantly, we use the optimal transport to compute pixel matching from a global perspective, thereby obtaining long-range correspondences. Quantitative results in Table 3 demonstrate that in the DreamGaussian-based model, introducing long-range correspondence results in a 2.8% improvement in CLIP score; in the TripoSR-based model, it results in a 4.5% improvement in CLIP score. However, after incorporating long-range correspondence, PSNR, SSIM, and LPIPS metrics are inferior compared to the local-range loss function. We conducted further qualitative studies (please see appendix) and found that without pixel matching, the layout cannot be aligned with the reference image. Additionally, when position distance and depth distance are introduced separately, there is no significant change in quantitative results, but there is a noticeable improvement in qualitative results, as illustrated in Figure 6.
Table 3: Comparison of different information used in appearance loss $L_{a}$ (Eq. 5). OT represents whether to use the optimal transport algorithm to obtain pixel matching.
(a) DreamGaussian-based REPARO.
Info in $L_{a}$ OT CLIP $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
RGB $\times$ 0.800 18.054 0.853 0.206
RGB 0.828 17.009 0.821 0.245
RGBXY 0.828 17.069 0.821 0.243
RGBDXY 0.833 17.391 0.827 0.232
(b) TripoSR-based REPARO.
Info in $L_{a}$ OT CLIP $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
RGB $\times$ 0.774 17.918 0.868 0.216
RGB 0.819 17.837 0.865 0.215
RGBXY 0.823 17.847 0.865 0.213
RGBDXY 0.822 17.765 0.865 0.216
@@
5 Conclusion and Discussion
@@
In conclusion, we introduce REPARO, a comprehensive approach for generating compositional 3D scenes from single images. Our REPARO addresses the multi-object issue by decomposing the scene into individual objects, reconstructing their 3D meshes, and optimizing their layout through differentiable rendering. During differentiable rendering, we incorporate an optimal transport-based long-range appearance loss which considers RGBDXY information, and a high-level semantic loss which aligns the feature correspondence. Consequently, REPARO is able to obtain multi-object 3D assets with visual-spatial arrangement and contextual consistency. Nevertheless, our findings indicate a discrepancy between the quantitative and qualitative results in the evaluation of multi-objective 3D assets. Specifically, the quantitative results show improvement but the qualitative results exhibit only marginal enhancement. Hence, it is imperative to develop evaluation methods for multi-object 3D assets in future research.
Appendix A Appendix
Figure 7: Qualitative comparison for local-range loss and our loss.
In Table 3, the loss term using only RGB color information without pixel matching obtains better PSNR, SSIM, and LPIPS performance compared to the method with long-range pixel matching. We show some qualitative results in Figure 7. These results demonstrate that the layout is not aligned with the reference image when using the local range loss term. By contrast, the proposed long-range loss term can optimize the scene parameters and align the layout with the reference image.

### v2

@@ 4.2 Main results
User study. In addition to numerical metrics, we conducted a user study to compare our method with others, gathering responses from participants who assessed the realism of different 3D assets. Our approach received 61% approval, while Dreamgaussian received 19% approval, LRM received 12% approval, and Wonder 3D received 8% approval. The results consistently showed that our approach was favored over previous methods. For more detailed results, please refer to the supplementary material.
@@ 4.3 Ablation study
Figure 5: Qualitative comparison for different options of optimization parameters $\theta$.
@@
Figure 7: 3D scene editing based on the implementation of 2D image editing. Based on our REPARO framework, it is possible to implement four types of 3D editing guided by 2D image editing, namely adding, removing, position exchanging and instance replacement.
Ablation for different loss functions in differentiable rendering. We conducted ablation experiments on different loss components in Eq. 9, with the results presented in Table 6. The results indicate that incorporating the high-level semantic loss function $L_{s}$ can bring improvements in both SSIM and LPIPS. However, the enhancement in CLIP score is not as pronounced. Additionally, when only $L_{s}$ is used, there is a notable decrease in CLIP score. For instance, in the DreamGaussian-based model, the CLIP score decreased by $2.1\%$. This suggests that relying solely on feature supervision can reduce the effectiveness of layout alignment.
Ablation for different options of scene parameters $\theta$. Table 6 provides an ablation study on the impact of various parameter optimization settings within $\theta$ across both DreamGaussian-based and TripoSR-based REPARO models. Specifically, we compare combinations of translation ($t$), rotation ($r$), and scale ($s$) optimizations. For DreamGaussian-based REPARO, the combination of parameters $t$ and $s$ achieves the highest CLIP score (0.833), competitive PSNR (17.279), and the lowest LPIPS (0.234), indicating a high level of perceptual quality and fidelity. Similarly, for TripoSR-based REPARO, optimizing $t$ and $s$ yields a balance between perceptual similarity (CLIP 0.822) and structural fidelity, with the lowest LPIPS (0.216) and strong SSIM (0.865), aligning closely with the input image.
Notably, as shown in Figure 5, including the rotation parameter $r$ in the optimization tends to degrade alignment with the input image orientation, as shown by the rotation-inclusive configurations, which generally exhibit slight reductions in CLIP scores and increased LPIPS. As a result, we select the combination of $t$ and $s$ for the final configuration, as it achieves optimal perceptual alignment without sacrificing the fidelity to the original image orientation.
Ablation for long-range appearance loss term. As shown in Eq. 5, long-range appearance loss $L_{a}$ incorporates RGB color distance, position distance, and depth value distance. More importantly, we use the optimal transport to compute pixel matching from a global perspective, thereby obtaining long-range correspondences. Quantitative results in Table 6 demonstrate that in the DreamGaussian-based model, introducing long-range correspondence results in a 2.8% improvement in CLIP score; in the TripoSR-based model, it results in a 4.5% improvement in CLIP score. Additionally, when position distance and depth distance are introduced separately, there is no significant change in quantitative results, but there is a noticeable improvement in qualitative results, as illustrated in Figure 6.
Table 7: DreamGaussian-based ($\clubsuit$), TripoSR-based ($\spadesuit$), and GT-based ($\blacklozenge$) results.
Method $\text{CLIP}_{\text{s}}$ $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
REPARO 0.833 17.279 0.826 0.234
REPARO 0.822 17.751 0.865 0.216
REPARO 0.906 15.708 0.789 0.321
The impact of 3D asset quality. As shown in Tab. 7, the experimental results indicate that the process of generating individual assets in the first stage introduces significant errors. Therefore, we still need a model capable of generating high-quality 3D assets.
4.4 Extension: Using 2D editing to guide 3D editing
We further explored potential applications of our combinatorial generation framework. As depicted in the Figure 7, the entire framework relies on a single input image as guidance, enabling both the reconstruction of individual objects and layout alignment through differentiable rendering. Leveraging this capability, we integrated pre-existing image editing techniques to facilitate editing of the generated 3D scenes. Specifically, we demonstrate four types of editing operations: adding, removing, position exchange, and instance replacement. The 2D editing process is implemented using the AnyDoor [2] model. By comparing the 3D scenes before and after editing, it is evident that the REPARO framework exhibits robust spatial arrangement capabilities and maintains consistency for each assets.
5 Conclusion
@@
In conclusion, we introduce REPARO, a comprehensive approach for generating compositional 3D scenes from single images. Our REPARO addresses the multi-object issue by decomposing the scene into individual objects, reconstructing their 3D meshes, and optimizing their layout through differentiable rendering. During differentiable rendering, we incorporate an optimal transport-based long-range appearance loss which considers RGBDXY information, and a high-level semantic loss which aligns the feature correspondence. Consequently, REPARO is able to obtain multi-object 3D assets with visual-spatial arrangement and contextual consistency. One limitation of our method is that severe occlusion may lead to hallucinated parts inconsistent with the image, and in semantically implausible scenes, layout optimization tends to favor common-sense interpretations over the actual image. We hope our method can pave the way for image to 3D generation in multi-object scenes.